OpenStudy (anonymous):

Can anybody please help me show my work for these problems....I have the correct answer just no work to prove it... 1.Given f(x) = x2 + 2x, find f(a – 3) A: a^2-4a+3 2. If f(x0=x/x-3 and g(x)=2x-1, find (f-g)(x) A: -2x^2+8x-3/x-3 3. If f(x)=x^2+1 and g(x)=1/x, find [f*g](x) A:1/x^2 +1 4. Which equation best represents a line perpendicular to the graph of x+3y=9? A:y=3x-1 5. Which of the following describes the graphs of 2x+5y=9 and 10x=4y+18? A: perpendicular 6. Write the standard form of an equation of the line parallel to the graph of x-2y-6=0 and passing through A(

OpenStudy (anonymous):

-3,2) A: x-2y+7=0 Also with... 7. Find the value of k such that -2/3 is a zero of the function f(x)=4x+k/7 A: 8/3 8. Find the zero of f(x)=-2/3x-12 A:-18 9. Line k pass through A(-3,-5) and has a slope of -1/3. What is the standard form of the equation for line k? A:x+3y+18=0 10. Write an equation in slope-intercept form for the line passing through A(-2,1) and having a slope of .5 A:y=.5x+2 11.Write an equation in standard form for a line with an x-intercept of 2 and a y-intercept of 5. A:5x+2y-10=0 12. Write the standard form of the equation of the line parallel to the graph of 2y-6=0 and passing through B(4,-1) A: y+1=0 13. Write an equation in standard form of the line perpendicular to the graph of x-2y-6=0 and passing through A(-3,2) A: 2x+y+4=0 Please just show me how to get the answer....I have the correct answer just not the work to receive full credit. I would really appreciate it(:

OpenStudy (amistre64):

so in other words, someone gave you the answers and not the work, now you want someone else to do the work?

OpenStudy (anonymous):

No it was a study sheet from my teacher with the correct answers, and she just wants to check our understanding...but I'm having problems in precal and don't know how to solve correctly...

OpenStudy (amistre64):

well, this might not be the right site for you then. If you want, post a single question that you need help with; allow the others to work with you on it to a suitable solution; then post another question. otherwise, noone is going to want to get suspended for simply typing up all the work. And just so that you are aware of the OpenStudy policy: It is better to guide the user to an answer rather than doing all the work yourself. We discourage giving out answers without letting the asker at least take some part in the analysis - the aim of this site is to teach. Please refer to our code of conduct at http://openstudy.com/code-of-conduct for more information. Thank you.

OpenStudy (amistre64):

i can go thru the first question with you if you want

OpenStudy (anonymous):

Ok, sorry... I just thought that this was a site where we help each other out... I mean I know how to start off the problems and the answer..but my teacher want the work not the answer so I was confused... I have read the code of conduct I just didn't think I could get in trouble for asking for the help...and I definitely don't want to get kicked off the site so I'll just seek help else where. But thank you so much for letting me know and sorry for not following the rules....good bye

OpenStudy (amistre64):

you havent broken any rules ... yet :) i just wanted to make sure you were aware of how we wish this site to operate. and, youll simply get better responses by posting the questions one at a time. I can still help you with the first question if you want.

OpenStudy (anonymous):

Oh ok, thanks!! Please, I would really appreciate it!!

OpenStudy (amistre64):

.Given f(x) = x2 + 2x, find f(a – 3) the notation f(x) relates to f(a-3) simply by saying that x=a-3; does that make sense?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

f(x) = x^2 + 2x so lets replace all the x parts with (a-3) f(a-3) = (a-3)^2 + 2(a-3) and the rest is just expanding it out from there

OpenStudy (hba):

@amistre64 Can i jump in ?

OpenStudy (amistre64):

only if its useful :)

OpenStudy (hba):

ok.

OpenStudy (anonymous):

Oh ok, I get it!!

OpenStudy (amistre64):

good luck with the rest of them; go ahead and post a new question in the ask a question box and im sure others will help out :)

OpenStudy (hba):

You may use this formula (a-b)^2=a^2-2ab+b^2 to expand it.

OpenStudy (anonymous):

Ok thanks so much....I do have a question how do we report a user?

OpenStudy (hba):

Post it in the Os feedback group.

OpenStudy (amistre64):

at the bottom of each post is a small print; says report abuse. Click it and fill it out

OpenStudy (anonymous):

This site is for studying and he was asking very inappropriate questions

OpenStudy (anonymous):

Ok thank you both!!